OpenStudy (anonymous):

The equation of line CD is y = −2x − 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5).

OpenStudy (♪chibiterasu):

Using y = mx + b, what is the slope of this lie given?

OpenStudy (♪chibiterasu):

line* .-.

OpenStudy (♪chibiterasu):

Well since you won't respond, the slope is -2, and since parallel lines have the same slope the new line will also have a slope of 2. Point slope form looks like this: \[(y _1 - y _2) = m(x _1 - x_2)\] Where m is the slope. So now with the slope and points given we can determine that the equation of the line will be: \[(5 - y _2) =- 2(4 - x _2)\]

OpenStudy (♪chibiterasu):

Remove the 2 from y and x.